OpenStudy (anonymous):

Find the potential real zeros (roots) of the following equation. f(x) = 4x5 - 16x4 + 17x3 - 19x2 + 13x - 3 = 0

OpenStudy (anonymous):

do you still want help on this one?

OpenStudy (anonymous):

X1 = 1/2 X2 = 1/2 X3 = 3 X4 = i X5 = - i

OpenStudy (anonymous):

these are your roots :)

OpenStudy (anonymous):

sorry I took so long I stepped of for a min.

OpenStudy (anonymous):

??

OpenStudy (anonymous):

the answers are in the second post

OpenStudy (anonymous):

sorry im just having major shockwave problems,but thanks I have it

OpenStudy (anonymous):

np